OpenStudy (hatahateo):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 3, -13, and 5 + 4i

OpenStudy (hatahateo):

Answers: f(x) = x4 - 8x3 - 12x2 + 400x - 1599 f(x) = x4 - 200x2 + 800x - 1599 f(x) = x4 - 98x2 + 800x - 1599 f(x) = x4 - 8x3 + 12x2 - 400x + 1599

OpenStudy (hatahateo):

@welshfella

OpenStudy (hatahateo):

This one is really confusing me

OpenStudy (welshfella):

Ok Do you know that all complex roots occur as conjugate pairs?

OpenStudy (hatahateo):

Yes

OpenStudy (welshfella):

So the remaining root is 5 - 4i so we have f(x) = (x - 3)(x + 13)( x - (5+4i)) ( x - ( 5 - 4i))

OpenStudy (hatahateo):

Alright

OpenStudy (welshfella):

expanding that is a bit of a pain , I know!

OpenStudy (hatahateo):

Yeah definitely haha

OpenStudy (hatahateo):

So then, I'm confused at what to do next

OpenStudy (welshfella):

work out the 2 complex factors first

OpenStudy (hatahateo):

I got (x-3)(x+13)(x-5-4i)(x-5+4i) I think I did something wrong, that doesn't seem right..

OpenStudy (hatahateo):

I actually got 41 by working out the 2 roots

OpenStudy (welshfella):

you should be able to eliminate the imaginary parts

OpenStudy (welshfella):

yes so that makes it (x^2 + 41)

OpenStudy (hatahateo):

Alright

OpenStudy (welshfella):

- see what you mean though - the expansion doesnt look right though - Maybe I got it wrong....

OpenStudy (hatahateo):

This question was very confusing to me and didn't make much sense

OpenStudy (welshfella):

yea I am a bit confused myself lol I cant see anything wrong with my algebra though...

OpenStudy (hatahateo):

So what would be my best bet with this question, it is my last one haha

OpenStudy (welshfella):

well expand (x - 3)(x + 13)(x^2 + 41) and hopefully it wil give 1 of choices

OpenStudy (welshfella):

wait its not x^2 + 41 we must have a term in x in it!!

OpenStudy (welshfella):

lets do it on a scribble pad

OpenStudy (welshfella):

yea it expands to x^2 - 5x + 4xi - 5x + 25 - 20i -4xi + 20i - 16 i^2

OpenStudy (welshfella):

there you go You'll see that you lose all the i's when you simplify that.

OpenStudy (welshfella):

remember that i^2 = -1

OpenStudy (welshfella):

so - 16i^2 = -16 * -1 = + 16

OpenStudy (welshfella):

so can you simplify it?

OpenStudy (hatahateo):

Let me try real quick

OpenStudy (hatahateo):

Not sure if this is right but I got -224

OpenStudy (hatahateo):

Yeah that doesn't seem right

OpenStudy (welshfella):

??? you should end up with a trinomial something like x^2 - ax + 41

OpenStudy (hatahateo):

I'm not sure what I did..

OpenStudy (welshfella):

the terms ending in i cancel out and last term = 25 + 16 = 41 and youl'l have a term in x

OpenStudy (welshfella):

Just add like terms

OpenStudy (hatahateo):

Okay

OpenStudy (welshfella):

x^2 - 5x + 4xi - 5x + 25 - 20i -4xi + 20i - 16 i^2 = x^2 - 5x - 5x + 25 + 16

OpenStudy (welshfella):

4xi and 20i disappear

OpenStudy (hatahateo):

Okay, I get that now

OpenStudy (welshfella):

so now you have to expand (x - 3)(x + 13)(x^2 - 10x + 41)

OpenStudy (hatahateo):

yes, I did that up there and got the same trinomial

OpenStudy (welshfella):

note that last term will be -3 + 13 * 41 = -1599 so you can eliminate D

OpenStudy (hatahateo):

I just got that too

OpenStudy (hatahateo):

x^4-98x^2+800x-1599

OpenStudy (hatahateo):

C?

OpenStudy (welshfella):

= ( x^2 + 10x - 39)(x^2 - 10x + 41)

OpenStudy (welshfella):

yes thats it notice the terms in x^3 are + 10x^3 and - 10x^3 so that eliminates them

OpenStudy (hatahateo):

Thanks! It was right

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

these are very tedious to work out and its easy to make a mistake.

OpenStudy (welshfella):

I was too lazy to do the expansion!